# Time-Dependent Black Hole Lensing and the Unified Weak-to-Strong Deflection Framework

### v1

@@ Appendix C Near-critical expansion and SDL coefficients
We comment that the cancellation of $a^{\prime}(r_{\mathrm{ph}})$ in Eq. (100) is a special feature of Schwarzschild and the particular combination entering $b_{c}$. It reflects the fact that $b_{c}$ depends on $A$ through $r_{\mathrm{ph}}$ and the normalization $A(r_{\mathrm{ph}})$ only; at linear order these two contributions conspire to eliminate the derivative at $r_{\mathrm{ph}}$.
@@
Logarithmic and finite response: $a_{1}(\omega)$ and $b_{1}(\omega)$: We next relate the coefficients $a_{1}(\omega)$ and $b_{1}(\omega)$ in Eq. (39) to moments of the near-photon-sphere Born kernel. Let $\hat{\bm{e}}_{b}$ be the radial unit vector on the screen. Axisymmetry implies that, to leading order in $\Delta=b/b_{c}-1$, the deflection correction $\delta\bm{\alpha}$ is radial, so we write $\delta\bm{\alpha}=\delta\alpha_{\parallel},\hat{\bm{e}}_{b}$. Comparing Eq. (37) to the scalar linearization (41) gives
$a_{1}(\omega)=-\hat{\bm{e}}_{b}\,\cdot\,\bm{\mathcal{C}}(\omega),\qquad b_{1}(\omega)=\hat{\bm{e}}_{b}\,\cdot\,\bm{\mathcal{D}}(\omega),$ (103)
@@
where $\bm{\mathcal{C}}=(\mathcal{C}_{X},\mathcal{C}_{Y})$ and $\bm{\mathcal{D}}=(\mathcal{D}_{X},\mathcal{D}_{Y})$ are the complex vector amplitudes in Eq. (37). In turn, $\mathcal{C}_{A}$ and $\mathcal{D}_{A}$ follow from the near-photon-sphere reduction of the kernel $B.11$. Introducing the azimuth $\varphi$ during the whirl phase (Eq. (36)) and the delay $\tau(\varphi)$, we can write
$\mathcal{C}_{A}(\omega)=\int^{\varphi_{\max}}_{\varphi_{\min}}d\varphi\,W_{A}^{(0)}(\varphi),$
$\mathcal{D}_{A}(\omega)=\int^{\varphi_{\max}}_{\varphi_{\min}}d\varphi\,\left[W_{A}^{(1)}(\varphi)+i\omega\,U_{A}^{(1)}(\varphi)\right],$ (104)
@@
where the weights $W_{A}^{(0)}$, $W_{A}^{(1)}$, $U_{A}^{(1)}$ are smooth functions of the background quantities at $r\simeq r_{\mathrm{ph}}$ and of the RWZ amplitudes $\Psi_{Z}\,\partial_{r_{*}}\Psi_{Z}$ evaluated there (see Eqs. (77) and (94)). Their explicit algebraic forms are lengthy but completely fixed by Eqs. (68), (72)-(75), and (88). The limits $(\varphi_{\min},\varphi_{\max})$ scale as $|\ln\Delta|$ and generate the logarithm in Eq. (37); the finite parts define $b_{1}$ once the inner and outer contributions are combined as in Eq. (52).
@@
Eqs. (103)-(104) summarize the content of $a_{1}$ and $b_{1}$: both are localized, frequency-resolved moments of the Zerilli field and its first radial derivative at $r_{\mathrm{ph}}$, with $a_{1}$ capturing the universal logarithmic weight and $b_{1}$ the finite near-photon-sphere remainder after matching to the outer integral (Sec. V.1).
@@
Time-of-flight counterparts: The coordinate travel time for near-critical rays has the structure as Eq. (45),
@@
$+\tilde{b}(t_{o})+n\,T_{\mathrm{whirl}}(t_{o})+\ldots,$ (105)
@@
$\tilde{a}_{1}(\omega)=-\hat{\bm{t}}\,\cdot\,\bm{\mathcal{C}}^{(T)}(\omega),\qquad\tilde{b}_{1}(\omega)=\hat{\bm{t}}\,\cdot\,\bm{\mathcal{D}}^{(T)}(\omega),$
$\gamma_{1}(\omega)=\frac{\delta T_{\mathrm{whirl}}}{T_{\mathrm{whirl}}},$ (106)
where $\hat{\bm{t}}$ denotes projection onto the time leg of the tetrad and the superscript $T$ reminds us that the weights are those appropriate to the time functional. The same near-photon-sphere data ${\Psi_{Z}\,\partial_{r_{*}}\Psi_{Z}}_{r_{\mathrm{ph}}}$ enter, so the inter-image delay modulation (Eq. (48)) and the spacing/ring modulations (Eqs. (49)-(50)) are tied to the same localized kernel, as emphasized in Sec. V.2.
Consistency with the $1/\Delta$ pole and matching: Expanding the time-dependent law $36$ to linear order using $b_{c}(t_{o})=b_{c},[1+\varepsilon\,\mathrm{Re}{e^{-i\omega t_{o}}\beta_{1}(\omega)}]$ and $\bar{a}(t_{o})=\bar{a}+\varepsilon\,\mathrm{Re}{e^{-i\omega t_{o}}a_{1}(\omega)}$ reproduces Eq. (41) with the identifications
@@
(ii) $a_{1}(\omega),b_{1}(\omega)\ \text{ from (C.9)-(C.10)}.$ (107)
The $1/\Delta$ term $\bar{a},\beta_{1}/\Delta$ is fixed entirely by (102) and is independent of the arbitrary split radius used in Eq. (51); the logarithmic and finite pieces are determined by the inner kernel moments and are rendered unique by the matching cancellations between Eqs. (52)-(53). This is the precise content of the overlap relation (55): the near-critical asymptotics of the $1/b$ Born amplitude encodes the triplet $(\beta_{1},a_{1},b_{1})$.
Near criticality, all time-dependent SDL coefficients are controlled by local data at the photon sphere together with universal whirl-integrals of the Born kernel:
  • The critical-scale modulation is quasi-local:
    $\beta_{1}(\omega)=\frac{1}{4}\,H_{0}(r_{\mathrm{ph}};\omega),$ (108)
    equivalently $\delta b_{c}/b_{c}=(3/2),a(r_{\mathrm{ph}})$ with $a=-h_{tt}$, and $P_{2}(0)=-1/2$.
  • The logarithmic and finite coefficients $a_{1}(\omega)$ and $b_{1}(\omega)$ are given by kernel moments in Eq. (104) built from $\Psi_{Z}$ and $\partial_{r_{*}}\Psi_{Z}$ at $r_{\mathrm{ph}}$, projected radially according to Eq. (103).
  • The time-of-flight counterparts $\tilde{a}_{1},\tilde{b}_{1},\gamma_{1}$ arise from the same local ingredients via Eq. (106), explaining the phase-locked modulations of spacing and delays (Sec. IV.2) and their correlation with ring-radius oscillations (Sec. V.2).
These results complete the derivation of Eqs. (38)-(41) from the kernel formalism, making explicit how the Zerilli master field, evaluated at the photon sphere, controls all near-critical, time-dependent imaging phenomenology in our unified framework.

### v2

@@ Appendix C Near-critical expansion and SDL coefficients
We comment that the cancellation of $a^{\prime}(r_{\mathrm{ph}})$ in Eq. (C.7) is a special feature of Schwarzschild and the particular combination entering $b_{c}$. It reflects the fact that $b_{c}$ depends on $A$ through $r_{\mathrm{ph}}$ and the normalization $A(r_{\mathrm{ph}})$ only; at linear order these two contributions conspire to eliminate the derivative at $r_{\mathrm{ph}}$.
@@
Logarithmic and finite response: $a_{1}(\omega)$ and $b_{1}(\omega)$: We next relate the coefficients $a_{1}(\omega)$ and $b_{1}(\omega)$ in Eq. (IV.7) to moments of the near-photon-sphere Born kernel. Let $\hat{\bm{e}}_{b}$ be the radial unit vector on the screen. Axisymmetry implies that, to leading order in $\Delta=b/b_{c}-1$, the deflection correction $\delta\bm{\alpha}$ is radial, so we write $\delta\bm{\alpha}=\delta\alpha_{\parallel},\hat{\bm{e}}_{b}$. Comparing Eq. (IV.5) to the scalar linearization (IV.9) gives
$a_{1}(\omega)=-\hat{\bm{e}}_{b}\,\cdot\,\bm{\mathcal{C}}(\omega),\qquad b_{1}(\omega)=\hat{\bm{e}}_{b}\,\cdot\,\bm{\mathcal{D}}(\omega),$ (C.10)
@@
where $\bm{\mathcal{C}}=(\mathcal{C}_{X},\mathcal{C}_{Y})$ and $\bm{\mathcal{D}}=(\mathcal{D}_{X},\mathcal{D}_{Y})$ are the complex vector amplitudes in Eq. (IV.5). In turn, $\mathcal{C}_{A}$ and $\mathcal{D}_{A}$ follow from the near-photon-sphere reduction of the kernel $B.11$. Introducing the azimuth $\varphi$ during the whirl phase (Eq. (IV.4)) and the delay $\tau(\varphi)$, we can write
$\mathcal{C}_{A}(\omega)=\int^{\varphi_{\max}}_{\varphi_{\min}}d\varphi\,W_{A}^{(0)}(\varphi),$ (C.11)
$\mathcal{D}_{A}(\omega)=\int^{\varphi_{\max}}_{\varphi_{\min}}d\varphi\,\left[W_{A}^{(1)}(\varphi)+i\omega\,U_{A}^{(1)}(\varphi)\right],$ (C.12)
@@
where the weights $W_{A}^{(0)}$, $W_{A}^{(1)}$, $U_{A}^{(1)}$ are smooth functions of the background quantities at $r\simeq r_{\mathrm{ph}}$ and of the RWZ amplitudes $\Psi_{Z}\,\partial_{r_{*}}\Psi_{Z}$ evaluated there (see Eqs. (A.13) and (B.18)). Their explicit algebraic forms are lengthy but completely fixed by Eqs. (A.4), (A.8)-(A.11), and (B.11). The limits $(\varphi_{\min},\varphi_{\max})$ scale as $|\ln\Delta|$ and generate the logarithm in Eq. (IV.5); the finite parts define $b_{1}$ once the inner and outer contributions are combined as in Eq. (V.2).
@@
Eqs. (C.10)-(C.12) summarize the content of $a_{1}$ and $b_{1}$: both are localized, frequency-resolved moments of the Zerilli field and its first radial derivative at $r_{\mathrm{ph}}$, with $a_{1}$ capturing the universal logarithmic weight and $b_{1}$ the finite near-photon-sphere remainder after matching to the outer integral (Sec. V.1).
@@
Time-of-flight counterparts: The coordinate travel time for near-critical rays has the structure as Eq. (IV.15),
@@
$+\tilde{b}(t_{o})+n\,T_{\mathrm{whirl}}(t_{o})+\ldots,$ (C.13)
@@
$\tilde{a}_{1}(\omega)=-\hat{\bm{t}}\,\cdot\,\bm{\mathcal{C}}^{(T)}(\omega),\qquad\tilde{b}_{1}(\omega)=\hat{\bm{t}}\,\cdot\,\bm{\mathcal{D}}^{(T)}(\omega),$ (C.14)
$\gamma_{1}(\omega)=\frac{\delta T_{\mathrm{whirl}}}{T_{\mathrm{whirl}}},$ (C.15)
where $\hat{\bm{t}}$ denotes projection onto the time leg of the tetrad and the superscript $T$ reminds us that the weights are those appropriate to the time functional. The same near-photon-sphere data ${\Psi_{Z}\,\partial_{r_{*}}\Psi_{Z}}_{r_{\mathrm{ph}}}$ enter, so the inter-image delay modulation (Eq. (IV.18)) and the spacing/ring modulations (Eqs. (IV.19)-(IV.20)) are tied to the same localized kernel, as emphasized in Sec. V.2.
Consistency with the $1/\Delta$ pole and matching: Expanding the time-dependent law $36$ to linear order using $b_{c}(t_{o})=b_{c},[1+\varepsilon\,\mathrm{Re}{e^{-i\omega t_{o}}\beta_{1}(\omega)}]$ and $\bar{a}(t_{o})=\bar{a}+\varepsilon\,\mathrm{Re}{e^{-i\omega t_{o}}a_{1}(\omega)}$ reproduces Eq. (IV.9) with the identifications
@@
(ii) $a_{1}(\omega),b_{1}(\omega)\ \text{ from (C.9)-(C.10)}.$ (C.16)
The $1/\Delta$ term $\bar{a},\beta_{1}/\Delta$ is fixed entirely by (C.9) and is independent of the arbitrary split radius used in Eq. (V.1); the logarithmic and finite pieces are determined by the inner kernel moments and are rendered unique by the matching cancellations between Eqs. (V.2)-(V.3). This is the precise content of the overlap relation (V.5): the near-critical asymptotics of the $1/b$ Born amplitude encodes the triplet $(\beta_{1},a_{1},b_{1})$.
Near criticality, all time-dependent SDL coefficients are controlled by local data at the photon sphere together with universal whirl-integrals of the Born kernel: (i) The critical-scale modulation is quasi-local:
$\beta_{1}(\omega)=\frac{1}{4}\,H_{0}(r_{\mathrm{ph}};\omega),$ (C.17)
equivalently $\delta b_{c}/b_{c}=(3/2),a(r_{\mathrm{ph}})$ with $a=-h_{tt}$, and $P_{2}(0)=-1/2$; (ii) The logarithmic and finite coefficients $a_{1}(\omega)$ and $b_{1}(\omega)$ are given by kernel moments in Eq. (C.12) built from $\Psi_{Z}$ and $\partial_{r_{*}}\Psi_{Z}$ at $r_{\mathrm{ph}}$, projected radially according to Eq. (C.10); (iii) The time-of-flight counterparts $\tilde{a}_{1},\tilde{b}_{1},\gamma_{1}$ arise from the same local ingredients via Eq. (C.15), explaining the phase-locked modulations of spacing and delays (Sec. IV.2) and their correlation with ring-radius oscillations (Sec. V.2).
These results complete the derivation of Eqs. (IV.6)-(IV.9) from the kernel formalism, making explicit how the Zerilli master field, evaluated at the photon sphere, controls all near-critical, time-dependent imaging phenomenology in our unified framework.